# Achronal localization and representation of the causal logic from conserved current, application to massive scalar boson

### v1

@@ 1 Introduction
       Application to the massive scalar boson. The localizations of the massive scalar boson in Euclidean space (i.e., the Euclidean covariant positive operator valued normalized measures in $\mathbb{R}^{3}$, called POL in [6]) are determined by the integrals of a positive definite rotational invariant probability density $J_{0}$ over the regions of localization [6, sec. 6]. $J_{0}$ is the zeroth component of a covariant conserved current $\mathfrak{J}$ if and and only if the kernel is causal (6). This result is by [13]. Causal kernels have been studied in [13], [14], [6]. Under the almost irrelevant condition that the causal kernel is $C^{4}$ it is easy to verify that $\mathfrak{J}$ satisfies the assumption in (5) thus giving rise to a covariant achronal localization of the massive scalar boson.
@@
       Representation of the causal logic. Every (covariant) achronal localization determines uniquely a (covariant) representation of the causal logic and vice versa [6, (21)]. Obviously this one-to-one correspondence emphasizes further the relevance of achronal localization.
@@ 3 Divergence theorem on open sets with almost Lipschitz boundary
The following is inspired by Maggi [16, Remark 9.5, Theorem 9.6]. As to the notation, for $x\in\mathbb{R}^{n}$, $z\in\mathbb{R}^{n-1}$, $s>0$ put $x^{\prime}:=(x_{1},\dots,x_{n-1})$ and $x=(x^{\prime},x_{n})$, and let $B_{s}(x):=\{y\in\mathbb{R}^{n}:|y-x|<s\}$, $C(x,s):=\{y\in\mathbb{R}^{n}:|y^{\prime}-x^{\prime}|<s,|y_{n}-x_{n}|<s\}$, $D(z,s):=\{y\in\mathbb{R}^{n-1}:|y-z|<s\}$. Given an outer measure $\omega$ on $\mathbb{R}^{n}$ and $X\subset\mathbb{R}^{n}$, the trace (restriction) of $\omega$ on $X$ is the outer measure $\omega|X$ on $\mathbb{R}^{n}$ given by $\omega|X(F):=\omega(F\cap X)$. $\mathcal{H}^{n-1}$ denotes the $n-1$-dimensional Hausdorff measure on $\mathbb{R}^{n}$.
@@
       $\mathcal{H}^{n-1}|C(x,s)\cap M=\mathcal{H}^{n-1}|\operatorname{graph}\,u|_{D}$ is a Radon measure due to the area formula ([16, Theorem 9.1]) . Moreover, one has $\mathcal{H}^{n-1}(\operatorname{graph}\,u|_{D(x^{\prime},s)\setminus G})=\int_{D(x^{\prime},s)\setminus G}|(\nabla u(z),-1)|\operatorname{d}z=0$. ∎
The following preparatory result (3) comprises by the case $M_{0}=\emptyset$ the divergence theorem on bounded open sets with Lipschitz boundary (see also e.g. [16, Remark 9.5] and [1, A8.8]).
@@
Proof. (a) Assume first $\varphi\in C_{c}^{1}(C(x,s))$. By (3) the surface integral is well-defined. Following step one of the proof of [16, Theorem 9.3] the result follows due to $C(x,s)\cap\partial E=C(x,s)\cap M$ by definition (3)(b). Note that the weak gradient $\nabla f_{\delta}$ equals $\frac{1}{2\delta}(-\nabla u(z),1)$ at $z\in G$.
       (b) Now the vector field $\nu_{E}$ is constructed. Applying the result in (a) to all $\varphi\in C_{c}^{1}(C(x,s)\cap C(\bar{x},\bar{s}))$ shows that $\nu$ and $\bar{\nu}$ coincide a.e. on their common domain. There are countably many $C(x_{m},s_{m})$ from the localization in (3), which cover $M$. (Indeed, $M_{0}$ is closed and hence $G_{\delta}$. Since $\partial E$ is compact it follows that $M=\partial E\setminus M_{0}$ is $\sigma$-compact.) Thus we may compose from the corresponding $\nu_{m}$ by means of [16, Theorem 9.1] an $\mathcal{H}^{n-1}$-a.e. determined unit vector field $\nu_{E}$ on $\partial E$ which locally coincides a.e. with $\nu$ from (3).
@@
Indeed, the first claim holds as $\mathcal{H}^{n-1}(\partial E\cap A_{\delta})<\infty$ and $A_{\delta}\downarrow_{\delta}M_{0}$. The very definition of the Minkowski content [12, 3.2.27] implies the second claim.
@@
       According to [16, Theorem 8.1], $\mathcal{H}^{n-1}(D)<\infty$ if $D$ is the image under an injective Lipschitz function on $\mathbb{R}^{n-1}$ in $\mathbb{R}^{n}$ of a Lebesgue measurable set $C$ of finite Lebesgue measure. If $C$ is a Lebesgue null set then $\mathcal{H}^{n-1}(D)=0$.
@@
       According to [12, Theorem 3.2.39], $\mathcal{H}^{n-1}(D)=\mathcal{M}^{n-1}(D)=0$ holds, if $D\subset\mathbb{R}^{n}$ is closed and if there is a Lipschitz function $h:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ mapping some bounded $C\subset\mathbb{R}^{n-1}$ onto $D$ or, equivalently, if $h$ is locally Lipschitz defined on $\overline{C}$ with $h(C)=D$. The latter holds because $\overline{C}$ is compact.
@@ 4 Flux passing through a maximal achronal set
The result (4) essentially is an extension of [11, Proposition 37] to spacelike Cauchy surfaces. Note that [11, Proposition 37] is equally valid for a $C^{1}$ vector field in place of a smooth one. General spacetime theories study Cauchy surfaces. These are the sets which meet every inextendible timelike smooth curve exactly once [19, Chapter 14, Definition 28]. By [19, Chapter 14, Lemma 29] they meet even all inextendible causal smooth curves. According to [6, (9) Remark, (36) Theorem] a spacelike Cauchy surface is just a causal base, i.e., it is a spacelike [^2] A set is called spacelike according to sec. 2. set which intersects every causal line. It equals the graph of the corresponding $1$-Lipschitz function.
@@
       (b) The claim is that $\mathfrak{h}$ is locally Lipschitz. By [11, Proposition 37], $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$, $\Phi(s,x):=\gamma_{x}(s)$ is a diffeomorphism. Write $\Phi=:(\varphi,\phi)$.
@@
       (d) Let $C\subset\mathbb{R}^{3}$ be bounded Borel. Then $\mathcal{H}^{3}(\{0\}\times C)<\infty$, $\mathcal{H}^{3}(\operatorname{graph}\tau|_{C})<\infty$ by [16, Theorem 8.1].
@@
       Recall the diffeomorphism $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$. Hence $f:=\mathbb{R}\times]-R/2,R/2[\times]-R/2,R/2[\to\mathbb{R}^{4}$, $f(s,x_{1},x_{2}):=\Phi\big{(}s;x_{1},x_{2},\sqrt{R^{2}-x_{1}^{2}-x_{2}^{2}}\,\big{)}$ is injective, $C^{1}$ whence Lipschitz on compact sets, and $f([-S,S]\times[-R/3,R/3]\times[-R/3,R/3])=D^{\prime}$. The claim holds by [16, Theorem 8.1].
       (f) Let $C\subset\mathbb{R}^{3}$ be a compact Lebesgue null set. Then $\mathcal{H}^{3}(\{0\}\times C)=0$ and $\mathcal{H}^{3}(\mathfrak{h}(C))=0$ hold by [16, Theorem 8.1]. Indeed, the former is obvious, the latter holds true as $\mathfrak{h}$ is Lipschitz on $C$ being compact.
       (g) Now let $\eta\in C_{c}^{1}(\mathbb{R}^{4})$ with $0\leq\eta\leq 1$ and $\eta|_{\overline{E}_{R}}=1$. Put $\phi:=\eta\mathfrak{v}$. Then (3) yields $0=\int_{M}\mathfrak{v}\,\nu_{E_{R}}\operatorname{d}\mathcal{H}^{3}$ as $\mathfrak{v}$ satisfies the continuity equation. Moreover the integration over $L_{R}$ yields 00 as the integrand is 00. So using [16, Theorem 9.1] it follows $\int_{\varpi(\mathfrak{h}(B_{R}))}\big{(}v_{0}(\tau(x),x)-v(\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}x^{3}=\int_{B_{R}}v_{0}(0,x)\operatorname{d}x^{3}$.
       (h) Finally, as explained in [11, Proposition 37], the condition $\sigma(x)>0$ for $|x|<R$ in (c) can be removed and the limit $R\to\infty$ carried out. One ends up with (4.2).∎
@@ 5 Covariant achronal localization out of covariant conserved current
       There exist AL with the localization operators $T(\Delta)$ being orthogonal projections [6, (22) Theorem]. In this case the localization operators commute. However, a quantum mechanical system localized by a projection valued AL necessarily does not have a semi-bounded energy operator. This no-go result following from Hegerfeldt ’s well-known theorem regards a first type of Einstein causality requirement whose modern generalized reformulation [4, 6] we shall present below.
@@
In [11], (CC) condition was proved true in a special case where the sets $\Delta$ belong to smooth Cauchy surfaces and in [6], the final result has been established that an AL satisfies (CC) in full. As argued in [6], (CC) even necessitate achronal localization. In fact spacelike localization is not sufficient since (CC) induces the localization in achronal hyperplanes. This fact is reported in [6] and studied in detail in [7]. Hence we consider the study and the explicit construction of the achronal localization of the massive scalar boson to be of utmost relevance.
@@
As the integrand is nonnegative, $\pi_{\phi,\Lambda}$ is a $\sigma$-additive measure. By (a) and (4.3), $\pi_{\phi,\Lambda}(\Lambda)=||\phi||^{2}$. Due to (5), (5) $\phi\mapsto\pi_{\phi,\Lambda}(\Delta)$ is the quadratic form of a bounded Hermitian sesquilinear form on $\mathcal{D}$. Therefore by [11, Lemma 48] there is a bounded operator $T(\Delta)$, $0\leq T(\Delta)\leq I$ with $\pi_{\phi,\Lambda}(\Delta)=\langle\phi,T_{\Lambda}(\Delta)\phi\rangle$.
@@
       So $T_{\Lambda}$ is weakly $\sigma$-additive, which by [20, Theorem 4.28] implies the strong $\sigma$-additivity.
@@ 6 Covariant achronal localizations of the massive scalar boson
       To be explicit in the following we use the momentum representation with $L^{2}(\mathbb{R}^{3})$ being the space of states. By [5, (6.1), (11) Theorem] one has
@@
where $g:[m^{2},\infty[\to\mathbb{R}$ is continuous with $g(m^{2})=1$ such that the zeroth component $K_{0}$ of $\mathfrak{K}$ is a positive definite kernel on $\mathbb{R}^{3}$ (see also [5, (55) Corollary]).
@@
We turn to the family of localizations of the massive scalar boson obtained out of its stress energy tensor [17]. For a thorough treatment see [11, sec. 6]. The family is indexed by the normalized future-directed timelike fourvectors $\mathfrak{n}$, i.e. $\mathfrak{n}^{\cdot 2}=1$, $n_{0}>0$. The related currents read still in the momentum representation with $\phi\in C_{c}^{\infty}(\mathbb{R}^{3})$
@@
for all $\mathfrak{n}$, $g=(\mathfrak{a},A)$, $\mathfrak{x}$. In addition, for every $\mathfrak{n}$, $\phi$, one has $J_{\mathfrak{n},0}(\phi,;0,\cdot)\geq 0$ and $\int J_{\mathfrak{n},0}(\phi,;0,x)\operatorname{d}^{3}x=||\phi||^{2}$ as shown in [11, (64), Theorem 54]. Lemma (5), (5) hold for every $\mathfrak{J}_{\mathfrak{n}}$ by the same proofs due to (6.5). Finally, (5)(b) holds for every $J_{\mathfrak{n},0}$ by a proof analogous to that of (6). In summary, (5) applies to $\mathfrak{J}_{\mathfrak{n}}$. It follows

### v2

@@ 1 Introduction
       Application to the massive scalar boson. The localizations of the massive scalar boson in Euclidean space (i.e., the Euclidean covariant positive operator valued normalized measures in $\mathbb{R}^{3}$, called POL in [7]) are determined by the integrals of a positive definite rotational invariant probability density $J_{0}$ over the regions of localization [7, sec. 6]. $J_{0}$ is the zeroth component of a covariant conserved current $\mathfrak{J}$ if and and only if the kernel is causal (6). This result is by [14]. Causal kernels have been studied in [14], [15], [7]. Under the almost irrelevant condition that the causal kernel is $C^{4}$ it is easy to verify that $\mathfrak{J}$ satisfies the assumption in (5) thus giving rise to a covariant achronal localization of the massive scalar boson.
@@
       Representation of the causal logic. Every (covariant) achronal localization determines uniquely a (covariant) representation of the causal logic and vice versa [7, (21)]. Obviously this one-to-one correspondence emphasizes further the relevance of achronal localization.
@@ 3 Divergence theorem on open sets with almost Lipschitz boundary
The following is inspired by Maggi [17, Remark 9.5, Theorem 9.6]. As to the notation, for $x\in\mathbb{R}^{n}$, $z\in\mathbb{R}^{n-1}$, $s>0$ put $x^{\prime}:=(x_{1},\dots,x_{n-1})$ and $x=(x^{\prime},x_{n})$, and let $B_{s}(x):=\{y\in\mathbb{R}^{n}:|y-x|<s\}$, $C(x,s):=\{y\in\mathbb{R}^{n}:|y^{\prime}-x^{\prime}|<s,|y_{n}-x_{n}|<s\}$, $D(z,s):=\{y\in\mathbb{R}^{n-1}:|y-z|<s\}$. Given an outer measure $\omega$ on $\mathbb{R}^{n}$ and $X\subset\mathbb{R}^{n}$, the trace (restriction) of $\omega$ on $X$ is the outer measure $\omega|X$ on $\mathbb{R}^{n}$ given by $\omega|X(F):=\omega(F\cap X)$. $\mathcal{H}^{n-1}$ denotes the $n-1$-dimensional Hausdorff measure on $\mathbb{R}^{n}$.
@@
       $\mathcal{H}^{n-1}|C(x,s)\cap M=\mathcal{H}^{n-1}|\operatorname{graph}\,u|_{D}$ is a Radon measure due to the area formula ([17, Theorem 9.1]) . Moreover, one has $\mathcal{H}^{n-1}(\operatorname{graph}\,u|_{D(x^{\prime},s)\setminus G})=\int_{D(x^{\prime},s)\setminus G}|(\nabla u(z),-1)|\operatorname{d}z=0$. ∎
The following preparatory result (3) comprises by the case $M_{0}=\emptyset$ the divergence theorem on bounded open sets with Lipschitz boundary (see also e.g. [17, Remark 9.5] and [1, A8.8]).
@@
Proof. (a) Assume first $\varphi\in C_{c}^{1}(C(x,s))$. By (3) the surface integral is well-defined. Following step one of the proof of [17, Theorem 9.3] the result follows due to $C(x,s)\cap\partial E=C(x,s)\cap M$ by definition (3)(b). Note that the weak gradient $\nabla f_{\delta}$ equals $\frac{1}{2\delta}(-\nabla u(z),1)$ at $z\in G$.
       (b) Now the vector field $\nu_{E}$ is constructed. Applying the result in (a) to all $\varphi\in C_{c}^{1}(C(x,s)\cap C(\bar{x},\bar{s}))$ shows that $\nu$ and $\bar{\nu}$ coincide a.e. on their common domain. There are countably many $C(x_{m},s_{m})$ from the localization in (3), which cover $M$. (Indeed, $M_{0}$ is closed and hence $G_{\delta}$. Since $\partial E$ is compact it follows that $M=\partial E\setminus M_{0}$ is $\sigma$-compact.) Thus we may compose from the corresponding $\nu_{m}$ by means of [17, Theorem 9.1] an $\mathcal{H}^{n-1}$-a.e. determined unit vector field $\nu_{E}$ on $\partial E$ which locally coincides a.e. with $\nu$ from (3).
@@
Indeed, the first claim holds as $\mathcal{H}^{n-1}(\partial E\cap A_{\delta})<\infty$ and $A_{\delta}\downarrow_{\delta}M_{0}$. The very definition of the Minkowski content [13, 3.2.27] implies the second claim.
@@
       According to [17, Theorem 8.1], $\mathcal{H}^{n-1}(D)<\infty$ if $D$ is the image under an injective Lipschitz function on $\mathbb{R}^{n-1}$ in $\mathbb{R}^{n}$ of a Lebesgue measurable set $C$ of finite Lebesgue measure. If $C$ is a Lebesgue null set then $\mathcal{H}^{n-1}(D)=0$.
@@
       According to [13, Theorem 3.2.39], $\mathcal{H}^{n-1}(D)=\mathcal{M}^{n-1}(D)=0$ holds, if $D\subset\mathbb{R}^{n}$ is closed and if there is a Lipschitz function $h:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ mapping some bounded $C\subset\mathbb{R}^{n-1}$ onto $D$ or, equivalently, if $h$ is locally Lipschitz defined on $\overline{C}$ with $h(C)=D$. The latter holds because $\overline{C}$ is compact.
@@ 4 Flux passing through a maximal achronal set
The result (4) essentially is an extension of [12, Proposition 37] to spacelike Cauchy surfaces. Note that [12, Proposition 37] is equally valid for a $C^{1}$ vector field in place of a smooth one. General spacetime theories study Cauchy surfaces. These are the sets which meet every inextendible timelike smooth curve exactly once [20, Chapter 14, Definition 28]. By [20, Chapter 14, Lemma 29] they meet even all inextendible causal smooth curves. According to [7, (9) Remark, (36) Theorem] a spacelike Cauchy surface is just a causal base, i.e., it is a spacelike [^2] A set is called spacelike according to sec. 2. set which intersects every causal line. It equals the graph of the corresponding $1$-Lipschitz function.
@@
       (b) The claim is that $\mathfrak{h}$ is locally Lipschitz. By [12, Proposition 37], $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$, $\Phi(s,x):=\gamma_{x}(s)$ is a diffeomorphism. Write $\Phi=:(\varphi,\phi)$.
@@
       (d) Let $C\subset\mathbb{R}^{3}$ be bounded Borel. Then $\mathcal{H}^{3}(\{0\}\times C)<\infty$, $\mathcal{H}^{3}(\operatorname{graph}\tau|_{C})<\infty$ by [17, Theorem 8.1].
@@
       Recall the diffeomorphism $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$. Hence $f:=\mathbb{R}\times]-R/2,R/2[\times]-R/2,R/2[\to\mathbb{R}^{4}$, $f(s,x_{1},x_{2}):=\Phi\big{(}s;x_{1},x_{2},\sqrt{R^{2}-x_{1}^{2}-x_{2}^{2}}\,\big{)}$ is injective, $C^{1}$ whence Lipschitz on compact sets, and $f([-S,S]\times[-R/3,R/3]\times[-R/3,R/3])=D^{\prime}$. The claim holds by [17, Theorem 8.1].
       (f) Let $C\subset\mathbb{R}^{3}$ be a compact Lebesgue null set. Then $\mathcal{H}^{3}(\{0\}\times C)=0$ and $\mathcal{H}^{3}(\mathfrak{h}(C))=0$ hold by [17, Theorem 8.1]. Indeed, the former is obvious, the latter holds true as $\mathfrak{h}$ is Lipschitz on $C$ being compact.
       (g) Now let $\eta\in C_{c}^{1}(\mathbb{R}^{4})$ with $0\leq\eta\leq 1$ and $\eta|_{\overline{E}_{R}}=1$. Put $\phi:=\eta\mathfrak{v}$. Then (3) yields $0=\int_{M}\mathfrak{v}\,\nu_{E_{R}}\operatorname{d}\mathcal{H}^{3}$ as $\mathfrak{v}$ satisfies the continuity equation. Moreover the integration over $L_{R}$ yields 00 as the integrand is 00. So using [17, Theorem 9.1] it follows $\int_{\varpi(\mathfrak{h}(B_{R}))}\big{(}v_{0}(\tau(x),x)-v(\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}x^{3}=\int_{B_{R}}v_{0}(0,x)\operatorname{d}x^{3}$.
       (h) Finally, as explained in [12, Proposition 37], the condition $\sigma(x)>0$ for $|x|<R$ in (c) can be removed and the limit $R\to\infty$ carried out. One ends up with (4.2).∎
@@ 5 Covariant achronal localization out of covariant conserved current
       There exist AL with the localization operators $T(\Delta)$ being orthogonal projections [7, (22) Theorem]. In this case the localization operators commute. However, a quantum mechanical system localized by a projection valued AL necessarily does not have a semi-bounded energy operator. This no-go result following from Hegerfeldt ’s well-known theorem regards a first type of Einstein causality requirement whose modern generalized reformulation [5, 7] we shall present below.
@@
In [12], condition CC was proved true in a special case where the sets $\Delta$ belong to smooth Cauchy surfaces and in [7], the final result has been established that an AL satisfies CC in full. As argued in [7], CC even necessitate achronal localization. In fact spacelike localization is not sufficient since CC induces the localization in achronal hyperplanes. This fact is reported in [7] and studied in detail in [8]. Hence we consider the study and the explicit construction of the achronal localization of the massive scalar boson to be of utmost relevance.
@@
As the integrand is nonnegative, $\pi_{\phi,\Lambda}$ is a $\sigma$-additive measure. By (a) and (4.3), $\pi_{\phi,\Lambda}(\Lambda)=||\phi||^{2}$. Due to (5), (5) $\phi\mapsto\pi_{\phi,\Lambda}(\Delta)$ is the quadratic form of a bounded Hermitian sesquilinear form on $\mathcal{D}$. Therefore by [12, Lemma 48] there is a bounded operator $T(\Delta)$, $0\leq T(\Delta)\leq I$ with $\pi_{\phi,\Lambda}(\Delta)=\langle\phi,T_{\Lambda}(\Delta)\phi\rangle$.
@@
       So $T_{\Lambda}$ is weakly $\sigma$-additive, which by [21, Theorem 4.28] implies the strong $\sigma$-additivity.
@@ 6 Covariant achronal localizations of the massive scalar boson
       To be explicit in the following we use the momentum representation with $L^{2}(\mathbb{R}^{3})$ being the space of states. By [6, (6.1), (11) Theorem] one has
@@
where $g:[m^{2},\infty[\to\mathbb{R}$ is continuous with $g(m^{2})=1$ such that the zeroth component $K_{0}$ of $\mathfrak{K}$ is a positive definite kernel on $\mathbb{R}^{3}$ (see also [6, (55) Corollary]).
@@
We turn to the family of localizations of the massive scalar boson obtained out of its stress energy tensor [18]. For a thorough treatment see [12, sec. 6]. The family is indexed by the normalized future-directed timelike fourvectors $\mathfrak{n}$, i.e. $\mathfrak{n}^{\cdot 2}=1$, $n_{0}>0$. The related currents read still in the momentum representation with $\phi\in C_{c}^{\infty}(\mathbb{R}^{3})$
@@
for all $\mathfrak{n}$, $g=(\mathfrak{a},A)$, $\mathfrak{x}$. In addition, for every $\mathfrak{n}$, $\phi$, one has $J_{\mathfrak{n},0}(\phi,;0,\cdot)\geq 0$ and $\int J_{\mathfrak{n},0}(\phi,;0,x)\operatorname{d}^{3}x=||\phi||^{2}$ as shown in [12, (64), Theorem 54]. Lemma (5), (5) hold for every $\mathfrak{J}_{\mathfrak{n}}$ by the same proofs due to (6.5). Finally, (5)(b) holds for every $J_{\mathfrak{n},0}$ by a proof analogous to that of (6). In summary, (5) applies to $\mathfrak{J}_{\mathfrak{n}}$. It follows